what is the slope of x+7y=14
Found 2 solutions by chashmish, Alan3354:
Answer by chashmish(8)

x+7y=14
u bring 7 to the other side.
x+y= 14/7
x+y= 2
so obviously if 2 numbers =2
each one of them is 1

Answer by Alan3354(69443)

Solve for y:
x+7y=14
7y = -x + 14
y = (-1/7)x + 2
slope, m = -1/7
------------------
I have no clue what chashmish is doing.